OpenStudy (anonymous):

Can you help me with this? just guide me through how to do it......? A manufacturer of lawn mowers uses the function A(n)=(132n+75250)/(n) to model the average cost per lawn mower, in dollars, where n is the number of lawn mowers produced. a. Write the function that represents the horizontal asymptote. b. Describe the meaning of the horizontal asymptote in the context of the problem. c. What is the minimum number of lawn mowers that must be manufactured to bring the average cost per lawn mower down to $199? Round appropriately and state the answer in a complete sentence.

OpenStudy (phi):

\[ A(n)= \frac{132n+75250}{n} = 132+ \frac{75250}{n} \] as n gets bigger and bigger, the fraction 75250/*very big number* approaches zero.... and A(n) approaches 132 in other words, A(n) for large n approaches 132 on the x-y graph, you would say the x= 132